OpenStudy (anonymous):

what is the solution of the following system of equation? 2x-6=2y 3-2x=y

OpenStudy (anonymous):

What you need to do is sub in (3-2x) into the first equation for y then solve that for x and put your answer for x in to the second equation and solve for y. Then you would have an x and y value. 2x - 6=2(3-2x) 2x - 6=6-4x 2x+4x=6+6 6x=12 x=2 3-2x=y 3-2(2)=y -1=y